# RadarDistill: Boosting Radar-based Object Detection Performance via Knowledge Distillation from LiDAR Features

### 08 Mar 2024 v1

@@ 3.4 Proposal-based Feature Distillation
PFD employs Proposal-level Feature Matching to reduce the discrepancies between the high-level features. It utilizes predictions generated from the radar, guiding the radar network to create object features that closely mimic the high-level features of LiDAR, thereby achieving a more accurate alignment in properly identified proposals. Additionally, for erroneously identified proposals, such as false positives, PFD directs the model to suppress falsely activated features, thereby mirroring the lower activation levels typically seen in LiDAR features for such instances.
We employ the radar’s predicted heatmap to adaptively define the crucial region according to radar. Additionally, there is a significant scale disparity between the high-level BEV features generated by radar and LiDAR. PFD aligns the scales between the two modalities by making the student mimic the normalized values that are applied in the channel direction for both radar and LiDAR. DenseEnc perform input $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ by first downsampling to 1/16 of its size and then upsampling to 1/8 to generate the initial high-level BEV feature $F^{h^{(1)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Subsequently, $F^{h^{(1)}}_{rdr}$ and CMA’s output $F^{a^{(2)}}_{rdr}$ are concatenated and then fed through a convolution layer resulting in the second high-level BEV feature $F^{h^{(2)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Similar to AFD, PFD is applied twice to each of the multiple feature sets $F^{h^{(m)}}_{rdr}$.
The first step involves thresholding the predicted radar heatmap $H^{cls}_{rdr}$ and the GT heatmap $H^{cls}_{GT}$ to identify the regions of true positives, false positives, and false negatives. The regions are identified using the following operations
@@
where $H^{cls}_{rdr}$ represents the predicted radar classification heatmap, $H^{cls}_{GT}$ is the GT classification heatmap, and $\sigma$ is the threshold parameter set at 0.1.
TP and FN are considered important regions according to the foreground, while FP indicates regions where radar is mostly wrong. Consequently, TP and FN are categorized into one group, FP into another, leading to the generation of a scaling mask $M^{scale}$ for each defined area. The scaling mask is expressed as
$M^{scale}_{i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (16)
where $N$ denotes the number of pixels, while $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters for the respective regions. The alignment of feature scales between radar and LiDAR involves applying the channel-wise softmax to the high-level BEV features $F^{h}_{ldr}$ and $F^{h^{(m)}}_{rdr}$, which results in $S^{ldr}$ and $S^{h^{(m)}}$, i.e.,
$S^{mod}_{c,:,:}=\frac{{\text{e}xp}(\frac{F^{mod}_{c,:,:}}{\tau})}{\sum^{C}_{c^{*}=1}{\text{e}xp}(\frac{F^{mod}_{c^{*},:,:}}{\tau})},$ (17)
where $mod\in\{ldr,h^{(m)}\}$, $\tau$ is a hyper-parameter that controls the smoothing of the distribution and is used as 0.5. Similar to AFD, PFD is employed twice to train $S^{h^{(m)}}$ to imitate $S^{ldr}$ and is computed as
$L^{(m)}_{high}=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{scale}_{i,j}|S^{ldr}_{c,i,j}-S^{h^{(m)}}_{c,i,j}|.$ (18)
The final PFD loss value is derived from the average of the PFD loss values computed from multiple PFD. The final PFD loss is calculated from
$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}(L^{(m)}_{high}).$ (19)
3.5 Overall Loss Function
We freeze the teacher model that has been pre-trained on LiDAR point clouds and then train the student model end-to-end using radar point clouds. Finally, we compute the total loss as the weighted sum of the detection loss and the distillation losses as
@@ 3.5 Overall Loss Function
where $\gamma$ and $\delta$ are hyperparameters used to balance the scale of different distillation losses.
@@ 4.1 Experimental Setup
Datasets and metrics. We conduct the experiments on the nuScenes [1] dataset. This dataset contains 700, 150, 150 driving scenes for training, validation, and testing each captured with images from six cameras, points from five radars and one LiDAR. Our evaluation utilizes the official metrics from nuScenes, which include mean average precision (mAP) and the nuScenes detection score (NDS).
@@
Table 1: Performance comparison on nuScenes test set. The best performed metrics in the radar-only model are bolded. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves the state-of-the-art performance in all metrics.
@@
Table 2: Per-class comparison on nuScenes test set. The best performed metrics are bolded. ’C.V’, ’Ped.’, ’Motor.’, and ’T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
Implementation details. As a baseline model, we utilized PillarNet-18, which is PillarNet [23] with a ResNet-18 backbone. We used the Adam optimizer with a learning rate of 0.001, applying a one-cycle learning rate policy. We set weight decay at 0.01, and adjusting momentum between 0.85 and 0.95. We trained our baseline models on 4 NVIDIA RTX 3090 GPUs for 20 epochs, employing a total batch size of 16 with the Class-Balanced Grouping and Sampling (CBGS) strategy [44]. We applied random scene flipping along the axes, random rotation, scaling, translation and ground-truth box sampling as the data augmentation. We set our detection range on the $X$ and $Y$ axes to $[-54.0m,54.0m]$ and on the Z-axis to $[-3m,5m]$ , utilizing pillars of $(0.075m,0.075m)$. Our proposed model was trained for $40$ epochs and all other training procedures were applied identically to the baseline. We adopted the teacher’s weights to initialize the student model, following the inheriting strategy [12]. In Eq. (9), we set the hyper-parameters $\alpha$ to 3e-4 and $\beta$ to 5e-5, while in Eq. (16), the hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ were set to $5$ and $1$, respectively. Additionally, in Eq. (20), we set the parameter $\gamma$ and $\delta$ to $5$ and $25$.
4.2 Comparison with the State-of-the-Arts
We evaluated the performance of RadarDistill on nuScenes test set. In Table 1, our RadarDistill achieved top-ranking performance on the nuScenes radar-only benchmark, significantly outperforming other methods. Specifically, our model achieved a $20.5\%$ mAP and $43.7\%$ NDS, surpassing the previous state-of-the-art (SOTA) method, KPConvPillars [28], by a substantial margin of $+15.6\%$ in mAP and $+29.8\%$ in NDS.
As shown in Table 2, we compared the performance of our RadarDistill for each class with other camera-only and camera-radar fusion models. Our model outperformed the baseline model (PillarNet [23]), especially for the barrier and trailer classes, with gains of at least $+23\%$. Furthermore, our model demonstrated the best performance for the car and trailer classes even when compared with other camera-only models and camera-radar fusion model approaches. However, the pedestrian and bicycle classes showed only slight performance improvements, largely due to the inherent limitations of radar sensors, as these objects are smaller and have few radar points in their ground truth (GT) boxes.
@@ 4.3 Ablation Studies
We conducted several ablation studies using the nuScenes validation set. We trained our model on 1/7 of training set and used the complete validation set for evaluation to reduce the time spent on experiments.
Component Analysis. We present an ablation study to evaluate the impact of each component of our proposed model, as shown in Table 3. In #2, when CMA is applied to the baseline, there’s an overall enhancement in performance; yet, the improvements are relatively low. Comparing #2 with #3 and #4, where CMA is combined with AFD and PFD respectively, we see a significant performance gain. For #3, there is an increase in mAP by $+4.5\%$ and NDS by $+4.4\%$. In the case of #4, mAP improves by $+1.8\%$ and NDS by $+3.0\%$. Additionally, when all components are used together as in #6, we can see that it achieves the best performance across all metrics. Lastly, to understand the influence of CMA in the distillation process, #5 was tested without it. The results show a significant difference in performance, with an increase of $+4.2\%$ in mAP and $+4.0\%$ in NDS when CMA is included, demonstrating that the CMA structure performs a critical role in our model.
Comparative Analysis of Distillation Regions. We conducted experiments to analyze the importance of distillation regions during the knowledge distillation (KD) process.
Table 4 presents an ablation study conducted to evaluate the impact of selecting different regions when applying low-level feature distillation. When distillation is performed on the entire area without separate regions, we observe a performance improvement of $+4.2\%$ in Car AP, $+2.2\%$ in mAP, and $+2.9\%$ in NDS. Applying distillation to the gaussian heatmap areas centered around the GT boxes led to an increase in Car AP and mAP compared to the total, but a decrease in NDS. Dividing the distillation between foreground and background regions, compared to the total, resulted in an increase in both Car AP and mAP, but a slight decrease in NDS. Conversely, dividing the feature distillation by our proposed activation regions yielded the best performance across all metrics, with improvements of $+6.8\%$ in Car AP, $+2.3\%$ in mAP, and $+1.5\%$ in NDS compared to total.
Next, Table 5 presents an experiment comparing the performance between previous cross-modality KD methods and PFD’s approach in applying high-level feature distillation. Here, the first row is based on the application of AFD according to our proposed method. When distillation is performed on the entire area, we observe a slight decrease in Car AP and mAP by $-0.6\%$ and $-0.1\%$ respectively, while NDS shows a marginal increase of $0.1\%$, indicating a minimal change in performance. Even when comparing the gaussian heatmap areas centered around the GT boxes, only Car AP shows a slight increase $0.1\%$ with the rest being identical to the total. Furthermore, applying KD based on LiDAR prediction results also did not yield any notable performance improvements. In contrast, our method, which applies KD based on radar prediction boxes, showed an improvement of over $0.4\%$ across all metrics compared to the total, demonstrating the best performance.
Scale Alignment in PFD. We conducted experiments to examine the impact of aligning the scale between teacher and student in PFD process in Table 6. When applying PFD without normalization, we observe a substantial decrease of $-5.9\%$ in Car AP, $-2.4\%$ in mAP, and $-1.9\%$ in NDS compared to before applying PFD. Conversely, after implementing normalization and PFD, we observe an increase of $+0.3\%$ in Car AP, $+0.3\%$ in mAP, and $+1.0\%$ in NDS.
Apply radar-camera fusion model. Table 7 shows a comparison of the performance when using radar features enhanced by our method in the radar-camera fusion model. The enhanced radar and camera features were fused using the BEVFusion [20] method, with its detailed structure and training method provided in the Supplementary Materials. We observe a significant increase of $+1.8\%$ in Car AP, $+1.3\%$ in mAP, and $+1.1\%$ in NDS.
@@
1 5.4 27.3
2 6.4 29.3
3 10.9 33.7
4 8.2 32.3
5 7.0 30.7
6 11.2 34.7
Table 3: Ablation study of components in Radardistill. “CMA”: Cross-Modality Alignment. “AFD”: Activation-based Feature Distillation. “PFD”: Proposal-based Feature Distillation.
@@
W/O AFD 34.8 6.4 29.3
Total 39.0 8.6 32.2
GT 40.4 8.8 29.5
FG / BG 42.1 9.6 31.7
Activation 45.8 10.9 33.7
Table 4: Ablation study of distillation regions in AFD. “W/O AFD”: Applying only CMA without distillation. “FG/BG”: Foreground and background region.
@@
W/O PFD 45.8 10.9 33.7
Total 45.2 10.8 33.8
GT 45.9 10.8 33.8
LiDAR prediction 45.4 10.8 34.0
Radar prediction 46.1 11.2 34.7
Table 5: Ablation study of distillation regions in PFD. “W/O PFD”: Applying only CMA and AFD. “LiDAR prediction” : Using only LiDAR prediction results. “Radar prediction” : Using only Radar prediction results
@@
W/O PFD 45.8 10.9 33.7
W/O Norm 39.9 8.5 31.8
Norm 46.1 11.2 34.7
Table 6: Ablation study of scale alignment in PFD. “W/O PFD”: Applying only CMA and AFD. “W/O Norm”: Applying PFD without normalizing. “Norm”: Applying normalizing and PFD.
@@
Table 7: Ablation study of applying to a radar-camera fusion model. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves slight performance enhancements in all metrics.
5 Conclusion
In our study, we introduced a new radar-based 3D object detection method, RadarDistill, which leveraged LiDAR data via knowledge distillation for enhanced performance. Our approach successfully guided the radar encoding network to generate radar features that closely resemble the detailed and semantically rich features of LiDAR. CMA enhanced the student network’s expressive power beyond that of the teacher network, allowing the student network to learn the complex distribution of LiDAR features more effectively. Through Activation-aware Feature Matching and Proposal-level Selective Feature Matching, both AFD and PFD successfully reduced the discrepancy from LiDAR feature distributions, particularly in important regions within the intermediate features. Our experiments demonstrated that RadarDistill achieved the state-of-the-art performance in radar-only object detection and offered significant performance gains in camera-radar fusion scenarios. We leave the design of more sophisticated KD framework specifically tailored for camera-radar fusion for future work.

### 05 Apr 2024 v2

@@ 3.4 Proposal-based Feature Distillation
PFD employs Proposal-level Feature Matching to transfer knowledge from the high-level features of LiDAR to those of radar. This guidance helps the radar network generate object features that closely mimic the high-level features of LiDAR, within object proposals.
DenseEnc is applied to produce two high-level features $F^{(h_{1})}_{ldr}$ and $F^{(h_{2})}_{ldr}$ from the low-level features $F^{(l)}_{ldr}$ in the LiDAR branch. Similarly, DenseEnc also produces two high-level features $F^{(h_{1})}_{rdr}$ and $F^{(h_{2})}_{rdr}$ from $F^{(l_{2})}_{rdr}$ in the radar branch. DenseEnc generates the first high-level features $F^{(h_{1})}_{mod}$ using the Conv2D-BN-ReLU block followed by a 2D convolution block and a 2D transposed convolution. Here, the Conv2D-BN-ReLU block refers to a sequence of operations comprising 2D Convolution (Conv2D), Batch Normalization (BN), and Rectified Linear Unit (ReLU), while the 2D convolution block consists of six layers of Conv2D-BN-ReLU. For LiDAR branch, DenseEnc produces the second high-level LiDAR features $F^{(h_{2})}_{ldr}$ by concatenate $F_{ldr}^{(h_{1})}$ with $F_{ldr}^{(l)}$ and applying the 2D Convolution block again. DenseEnc also generates the second high-level radar features $F^{(h_{2})}_{rdr}$ by concatenate $F_{rdr}^{(h_{1})}$ with $F_{rdr}^{(l_{2})}$ and applying the 2D Convolution block in radar branch. Finally, the CenterHead is applied to both $F^{(h_{2})}_{ldr}$ and $F^{(h_{2})}_{rdr}$ to produce the classification heatmaps $H^{cls}_{ldr}$ and $H^{cls}_{rdr}$, respectively.
PFD conducts knowledge distillation solely within the key target regions of high-level BEV features. PFD identifies key target regions based on the predicted radar heatmap $H^{cls}_{rdr}$ and the ground truth heatmap $H^{cls}_{GT}$. The ground truth heatmap is generated by projecting the 3D centers of ground truth boxes onto the BEV space and applying a Gaussian kernel to represent object locations [44]. The identified key target regions are labeled as true positives (TP), false positives (FP), or false negatives (FN) according to
@@
where $\sigma$ represents the threshold parameter set to 0.1 in our setup. TP and FN regions encompass areas containing real objects, while FP regions correspond to mis-detected regions. Consequently, PFD concentrates on feature alignment in both TP and FN regions while suppressing radar features in FP regions. Due to the inherent imbalance in the proportions of these three regions, distinct weights are applied to the distillation loss terms associated with TP, FN, and FP regions. The proposal-dependent loss weights $W_{proposal}$ are determined depending on the area of each region
$W_{proposal,i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (14)
where $\lambda_{1}$ and $\lambda_{2}$ represent the balancing parameters for the respective regions, and $N_{TP}$, $N_{FN}$ and $N_{FP}$ denote the number of pixels within TP, FN, and FP regions, respectively.
Finally, the distillation loss defined for the $m$th high-level features is weighted by the loss weights $W_{proposal}$ as
$L^{(m)}_{high}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}W_{proposal,i,j}\left|S^{(h_{m})}_{ldr,c,i,j}-S^{(h_{m})}_{rdr,c,i,j}\right|,$ (15)
where
$S^{(h_{m})}_{mod,c,i,j}=\frac{\text{exp}(F^{(h_{m})}_{mod,c,i,j})}{\sum^{C}_{k=1}\text{exp}(F^{(h_{m})}_{mod,k,i,j})},$ (16)
where $mod\in\{rdr,ldr\}$. Both the radar and LiDAR features $F^{(h_{m})}_{rdr}$ and $F^{(h_{m})}_{ldr}$ go through normalization in the channel dimension, resulting in $S^{(h_{m})}_{rdr}$ and $S^{(h_{m})}_{ldr}$, respectively. This normalization step aims to align the magnitude scale of the high-level features between the radar and LiDAR branches. The final PFD loss $L_{PFD}$ is computed by averaging the distillation loss over $m=1,2$, i.e.,
$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}L^{(m)}_{high}.$ (17)
3.5 Loss Function
The total loss function used to train RadarDistill is obtained from
@@ 3.5 Loss Function
where $\gamma$ and $\delta$ are the regularization parameters for weighting the loss terms $L_{AFD}$ and $L_{PFD}$, respectively.
@@ 4.1 Experimental Setup
Datasets and metrics. We conduct the experiments on the nuScenes [1] dataset. This dataset contains 700, 150, 150 driving scenes for training, validation, and testing, respectively. The nuScenes dataset gathers radar data using three radar sensors at the vehicle’s front and corners, and two at the rear, providing 360-degree coverage. These sensors collectively capture data at a frequency of 13Hz, ensuring broad coverage up to 250 meters and offering velocity measurement accuracy within ±$\pm 0.1km/h$. Our evaluation utilizes the official metrics from nuScenes, which include mean average precision (mAP) and the nuScenes detection score (NDS).
@@
Table 1: Performance evaluation on nuScenes testset. The best performed metrics in the radar-only model are marked in bold. “*” denotes models we have reproduced without applying test time augmentation. Our model achieves the state-of-the-art performance in all metrics.
@@
Table 2: Performance evaluation per class on nuScenes testset. The best performed metrics are marked in bold. ‘C.V’, ‘Ped.’, ‘Motor.’, and ‘T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
Implementation details. As our baseline model, we employed PillarNet-18, which corresponds to PillarNet [23] with a ResNet-18 backbone. We utilized the Adam optimizer with a learning rate of 0.001, implementing a one-cycle learning rate policy. The weight decay was set to 0.01 with the momentum scheduled from 0.85 to 0.95. We trained the PillarNet-18 model for 20 epochs with a batch size of 16. We used the Class-Balanced Grouping and Sampling (CBGS) strategy [45] to mitigate class imbalance issue. Data augmentation techniques were applied, including random scene flipping along the $X$ and $Y$ axes, random rotation, scaling, translation, and ground-truth box sampling. Our detection range was set to $[-54m,54m]$ for both the $X$ and $Y$ axes, and $[-5m,3m]$ for the $Z$ axis. We used pillars of dimensions $(0.075m,0.075m)$.
Our proposed RadarDistill model was trained for 40 epochs. The rest of training setup follows the same training setup as the baseline model. We initialized the radar backbone network using the pre-trained LiDAR backbone network, following the Inheriting Strategy presented in [12]. The hyperparameters $\alpha$ and $\beta$ in (8) were set to $3\times 10^{-4}$ and $5\times 10^{-5}$, respectively. Additionally, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ in (14) were set to $5$ and $1$, respectively. The parameters $\gamma$ and $\delta$ in (18) were set to $5$ and $25$, respectively. The entire training was conducted on 4 NVIDIA RTX 3090 GPUs.
4.2 Performance Comparison
Table 1 presents the performance of RadarDistill on nuScenes testset. Our RadarDistill achieves significant performance improvements compared to other radar-based object detectors. Specifically, RadarDistill achieves a mAP of 20.5% and a NDS of 43.7%, outperforming the previous state-of-the-art (SOTA) method, KPConvPillars [29], by a considerable margin of +15.6% in mAP and +29.8% in NDS. Table 2 displays the AP performance of RadarDistill for each class. We compare our RadarDistill with other camera-only and camera-radar fusion-based object detectors. RadarDistill exhibits superior performance compared to the baseline, PillarNet-18 [23] across all classes. Notably, it achieves performance gains of +12.2% and +23% for the Car and Trailer classes, respectively. Furthermore, RadarDistill outperforms other camera-only and sensor fusion models in the Car class, which is surprising given the limited resolution of radar data compared to camera data.
@@ 4.3 Ablation Studies
We conducted ablation studies on the nuScenes validation set to assess the impact of each proposed idea. Our model was trained using $1/7$ of the training set and evaluated on the entire validation set.
Component Analysis. Table 3 illustrates the performance enhancements achieved by each component of our proposed model. Integrating CMA into the PillarNet-18 baseline results in a 2% improvement in NDS performance. With both CMA and AFD enabled, the NDS performance further improves by 4.4%. Activating PFD yields a 1.0% additional increase in NDS performance. We also tried disabling CMA while other components are enabled in RadarDistill. In this case, the NDS performance drops by 4%, highlighting the critical role of CMA in our RadarDistill model.
Proposed Distillation Method. We compare the proposed Distillation method with other well known KD methods. We consider the following methods
  • Baseline*: This baseline is constructed by integrating CMA into the PillarNet-18 baseline.
  • Baseline**: This baseline is constructed by applying CMA and AFD to the PillarNet-18 baseline.
  • Complete [22]: This method applies the fixed weight across entire areas of the low-level BEV features.
  • Gaussian [26]: This approach generates a 2D Gaussian mask based on the centers of GT boxes and assigns higher weighted distillation loss to areas closer to the foreground.
  • FG/BG [7]: This strategy separates the foreground and background using GT boxes, applying different weighted distillation losses for each.
Table 4 evaluates the performance of AFD in comparison with other KD methods. We applied AFD and these methods to the Baseline*, where the low-level BEV features are produced by CMA. We observe that the proposed AFD achieves significantly better performance than other KD methods. Particularly, it achieves the performance gains of 3.7% in AP in car category, 1.3% in mAP, and 2.0% in NDS over the FG/BG method.
Table 5 compares PFD with other KD methods when applied to high-level BEV features. We applied these methods to Baseline** where the high-level features are produced by CMA and AFD. We also confirm that the proposed PFD achieves higher detection performance gains than other KD methods. Specifically, it yields the performance improvements of 0.5% in AP for the car category, 0.2% in mAP, and 1.0% in NDS over the FG/BG method.
@@
5.4 27.3
6.4 29.3
10.9 33.7
11.2 34.7
7.0 30.7
Table 3: Ablation study for evaluating the contribution from each component of Radardistill on nuScenes validation set.
@@
Baseline* 34.8 6.4 29.3
Complete [22] 39.0 8.6 32.2
Gaussian [26] 40.4 8.8 29.5
FG/BG [7] 42.1 9.6 31.7
Our AFD 45.8 10.9 33.7
Table 4: Comparison of AFD with different KD methods evaluated on nuScenes validation set.
Impact of Scale Normalization in PFD. Next, we investigate the impact of scale normalization in PFD. Table 6 presents a comparison of performances with and without scale normalization. We observe that our scale normalization results in a significant improvement of 6.2% in AP (CAR category), 2.7% in mAP, and 2.9% in NDS. This underscores the substantial difference in magnitude scale between radar and LiDAR high-level features, and demonstrates that narrowing this gap significantly boosts the effect of knowledge distillation.
Comparison when applied to Radar-Camera Fusion. Next, we investigate whether the improved features generated by RadarDistill can also lead to performance enhancements in radar-camera fusion methods. Table 7 presents the detection performance when the proposed RadarDistill is intergrated into the radar-camera fusion method, BEVFusion [20]. Because BEVFusion was originally designed for LiDAR-camera fusion, we adapted its design for radar-camera fusion following [15]. Then, RadarDistill was simply incorporated to the radar encoding network in BEVFusion and the entire model was trained end to end. We note that RadarDistill yields performance improvements of 1.8% in Car AP, 1.3% in mAP, and 1.1% in NDS over the baseline BEVFusion model. We believe that further enhancements in performance could be achieved by applying more sophisticated fusion strategies, a direction we leave for future research.
@@
Baseline** 45.8 10.9 33.7
Complete [22] 45.2 10.8 33.8
Gaussian [26] 45.9 10.8 33.8
FG/BG [7] 45.6 11.0 33.7
Our PFD 46.1 11.2 34.7
Table 5: Comparison of PFD with different KD methods evaluated on nuScenes validation set.
@@
PFD w/o norm 39.9 8.5 31.8
PFD 46.1 11.2 34.7
Table 6: Ablation study for evaluating the impact of scale normalization used in PFD on nuScenes validation set. PFD w/o norm: PFD without scale normalization. PFD: PFD with scale normalization.
@@
Table 7: Ablation study for evaluating the impact of RadarDistill on radar-camera fusion models. “*” denotes models we have reproduced without applying test time augmentation.
5 Conclusions
In this study, we introduced RadarDistill, a novel radar-based 3D object detection method aimed at enhancing radar data features by leveraging information from LiDAR data through knowledge distillation. We proposed three effective techniques—CMA, AFD, and PFD—to make the challenging task of transferring knowledge from LiDAR to radar successful. Our approach emphasizes guiding the radar encoding network to generate features that closely resemble the semantically rich features of LiDAR. CMA densifies radar features, aiding the radar encoding network in learning the complex distribution of LiDAR features better. AFD and PFD target the reduction of discrepancies between radar and LiDAR features in both low-level and high-level BEV features. Our evaluation demonstrated that RadarDistill can yield significant performance improvements over both radar-based and radar-camera fusion-based baselines, establishing state-of-the-art performance in radar-only object detection.
6 Acknowledgement
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.2021-0-01343-004, Artificial Intelligence Graduate School Program (Seoul National University)] and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No.2020R1A2C2012146].